Ethyl alcohol has a boiling point of 78.0 degree C
Answer #1

first it change from gas to liquid ( phase change )

Q = Lvm = 879 * .567 = 498.4 kJ

now from 78 degree to -114 degree

Q = mcdT

Q = .567 * 2.43 * 192 K = 264.54 kJ

now again phase change from liquid to solid

Q = Lf*m

Q = 109*0.567 = 61.803 kJ

total = 498.4 kJ + 264.54 kJ + 61.803

total energy removed = 824.74 kJ